OpenStudy (anonymous):

How many atoms are in a sample of 68.7 grams copper (Cu)?

OpenStudy (anonymous):

63.5gm Cu contains 6.023 into 10^23 atoms

OpenStudy (anonymous):

wouldnt it be 6.51 into 10^23?

OpenStudy (anonymous):

No. \[6.023 \times 10^{23}\] is a constant called Avogadro's number

OpenStudy (anonymous):

ya,Avogadros number is 6.022 so you would put 6.022x10^23 + 63.5/1 mol?

OpenStudy (jfraser):

no. 1 mole of Cu atoms is 6.022x10^23 atoms, and that pile will weigh 63.5g. You have a pile of copper atoms that weighs more than 63.5g. The conversions have to go:\[68.7g \space Cu * \frac{1mol \space Cu \space atoms}{63.5g \space Cu} * \frac{6.022*10^{23} \space atoms}{1 \space mol \space Cu \space atoms}\]see how all the units except for the # of atoms cancel? The molar mass is the link between the units of mass and quantity.

OpenStudy (anonymous):

with the given mass, the number of moles of Cu in the sample can be obtained: m(Cu)=68.7g n=m/M n(Cu)=68.7 / 63.55 =1.081 moles 1 mole= 6.022x10^23 atoms so 1.081 moles = (6.022x10^23)(1.081) =6.51x10^23 atoms